# Online and Offline Space-Filling Input Design for Nonlinear System Identification: A Receding Horizon Control-Based Approach

### v1

@@ IV EVALUATION
With the space-filling metrics, see Eq. (13) and Eq. (14), the process regressor space $\underline{X}$ is examined. Although this space is generally not accessible in practice, in this artificial setting, it is very useful for evaluating the exploration of the process’s operating regions of interest.
@@ IV-A Theoretical Analysis
In Fig. 2, an illustrative example of the regressor space distribution $\underline{{\hat{X}}}$ approximated by the surrogate model during an exemplary optimization procedure is shown, along with its corresponding supporting dataset $\underline{{\Psi}}$.
Figure 2: Exemplary illustration of the datasets employed during optimization, i.e., the regressor space distribution $\underline{\hat{X}}$ approximated by the surrogate model, along with the supporting dataset $\underline{{\Psi}}$.
The uniformly distributed supporting dataset $\underline{\Psi}$ is generated using Sobol sequences [27]. To enhance visualization, the figure displays an illustrative case, where $N_{\Psi}=N=100$. However, in practice, it is recommended to choose the number of supporting points such that $N_{\Psi}\gg N$. By default, $N_{\Psi}=5N$. This resolution is sufficient to efficiently target gaps when optimizing $\underline{\hat{{X}}}$, allowing to create high-quality space-filling distributions.
@@
Figure 3: Illustrations of the regressor space distribution, $\underline{{X}}_{a}$, and the corresponding signal trajectory, $\underline{u}_{a}$, resulting from the proposed method employing the adaptive modeling approach in (a) and (c), as well as their counterparts using the LTI as fixed model, $\underline{{X}}_{f}$ and $\underline{u}_{f}$, in (b) and (d).
@@ IV-B Space-Filling Performance
$\underline{U}_{\mathrm{IDSFID}}$ Excitation generated using the IDSFID algorithm with hyperparameter values $\lambda\in(0,0.5)$, resulting in a spectrum ranging from predominantly static to rather transient excitation. Employs the fixed LTI, as well.
@@
TABLE I: Properties of the investigated excitation signal methods. For each method, 50 signals of size $N=50$ were generated.
@@
In Fig. 5, the radius of the largest ball ${R}$ is shown as a function of the dataset size $N$. Displayed are the median-performing signals of the investigations shown in Fig. 4. Initially, the performance of $\underline{u}_{a}$ and $\underline{u}_{f}$ is similar. However, as the dataset size increases, the proposed method employing the adaptive modeling approach demonstrates superiority. This can be attributed to the continuous updating of model parameters, resulting in a more refined model as the number of process observations grows. This enables the efficient targeting of gaps in the data distribution. The same investigations as in Fig. 5 were conducted using the JSD. However, since the results are similar, they are not explicitly presented here.
Figure 5: Progress of the radius of the largest ball ${R}$ in dependence of the dataset size for different excitation signal design methods.
@@ V Discussion and Outlook
It can be concluded that high-quality surrogate models produce excitation signals rich in valuable information. However, the general quality of such models depends on prior knowledge and/or the availability of high-quality data – both are typically lacking in input design methods. The proposed approach uses minimal prior knowledge and can be considered as ”less model-based” than input design methods that rely on intrinsic model properties, such as parameter estimate variance and uncertainty. Furthermore, previous findings suggest that even a coarse process model can yield favorable results when optimizing a data distribution-centered optimality criterion [16]. Nonetheless, a systematic investigation into the relationship between model quality and the resulting input signal quality remains essential and will be conducted in the future. Further future research will focus on:
  • A systematic comparison of the employed spaces, specifically the joint input-state space and the regressor space, will be conducted.
  • Assessing whether distributions other than space-filling designs may be beneficial under specific conditions, like, e.g., solely local nonlinearities.
  • Analyzing the computational complexity.

### v2

@@ A Hammerstein process is employed as example process in the evaluation, defined by:
The prediction horizon is set to $L=4T$, ensuring that the sequence currently optimized $\tilde{\underline{U}}(k)$ can access all operating regions of the process and thus effectively target gaps in the model input space distribution.
With the space-filling metrics, see  (11) and  (12), the process regressor space
$\underline{X}=\begin{bmatrix}u(1)&\cdots&u(N)\\ y(1)&\cdots&y(N)\end{bmatrix}^{T}$ (14)
is examined. Here, $y(j)\ \forall j\in\{1,2,\ldots,N\}$ are obtained from  (13). Although this space is generally not accessible in practice, in this artificial setting, it is very useful for evaluating the exploration of the process’s operating regions.
@@ IV-A Theoretical Analysis
[…]
Figure 3: Illustrations of the process regressor space distribution $\underline{{X}}_{a}$, see (14), and the corresponding signal trajectory $\underline{u}_{a}$, generated using the proposed method with adaptive modeling, are shown in (a) and (c). Their counterparts, obtained using the fixed LTI model, $\underline{{X}}_{f}$ and $\underline{u}_{f}$, are shown in (b) and (d).
@@ IV-B Space-Filling Performance
$\underline{U}_{\mathrm{IDSFID}}$ Excitation generated using the IDS-FID algorithm with hyperparameter values $\lambda\in(0,0.5)$, resulting in a spectrum ranging from predominantly static to rather transient excitation. Employs the fixed LTI, as well.
@@
TABLE I: Properties of the investigated excitation signal methods. For each method, 50 excitation signals of size $N=300$ were generated.
@@
Figure 5 shows how the radius of the largest ball ${R}(k)$ evolves as the dataset size increases. Displayed are the median-performing signals of the investigations shown in Fig. 4. Initially, the performance of $\underline{u}_{a}$ and $\underline{u}_{f}$ is similar. However, as the dataset size increases, the proposed method employing the adaptive modeling approach demonstrates superiority. This improvement is attributed to the continuous updating of model parameters, which refines the model as more process observations are collected, thereby enabling efficient targeting of gaps in the data distribution. The same investigations as in Fig. 5 were conducted using the JSD. However, since the results are similar, they are not explicitly presented here.
[…]
Figure 5: Progress of the radius of the largest ball ${R}$ as a function of the dataset size for different excitation signal design methods. The median-performing signals of the investigations shown in Fig. 4 are displayed.
@@ V Discussion and Outlook
The proposed method addresses the generalization error by inducing a space-filling distribution in a surrogate model’s input space, achieving superior performance over existing input signal design approaches. Future work will assess the method on more complex and realistic processes to further validate these results.
The results further highlight that the effectiveness of model-based input design is closely linked to the surrogate model’s quality. Although the approach is less dependent on intrinsic model properties, such as parameter estimate variance and uncertainty, than traditional methods, a systematic analysis of the interplay between surrogate model accuracy and the resulting input signal quality is required and will be pursued.
Further research will focus on:
  • Systematic comparison of the employed spaces, specifically the joint input-state space and the regressor space.
  • Assessing whether distributions other than space-filling designs may be beneficial under specific conditions, such as the presence of solely local nonlinearities.
  • Analyzing computational complexity, especially in the context of online execution, the number of supporting points, and the length of the prediction horizon.
  • Evaluating the impact of the number of supporting points and the length of the prediction horizon on the input signal quality.